# Re-Reading Improves Reasoning in Large Language Models

### v3

@@ Abstract
To enhance the reasoning capabilities of off-the-shelf Large Language Models (LLMs), we introduce a simple, yet general and effective prompting method, Re2, i.e., Re-Reading the question as input. Unlike most thought-eliciting prompting methods, such as Chain-of-Thought (CoT), which aim to elicit the reasoning process in the output, Re2 shifts the focus to the input by processing questions twice, thereby enhancing the understanding process. Consequently, Re2 demonstrates strong generality and compatibility with most thought-eliciting prompting methods, including CoT. Crucially, Re2 facilitates a "bidirectional" encoding in unidirectional decoder-only LLMs because the first pass could provide global information for the second pass. We begin with a preliminary empirical study as the foundation of Re2, illustrating its potential to enable "bidirectional" attention mechanisms. We then evaluate Re2 on extensive reasoning benchmarks across 14 datasets, spanning 112 experiments, to validate its effectiveness and generality. Our findings indicate that, with the exception of a few scenarios on vanilla ChatGPT, Re2 consistently enhances the reasoning performance of LLMs through a simple re-reading strategy. Further analyses reveal Re2’s adaptability, showing how it can be effectively integrated with different LLMs, thought-eliciting prompting, and ensemble strategies.111  Our code is available at https://github.com/Tebmer/Rereading-LLM-Reasoning/
@@
Xiaohan Xu1thanks:   This work was done during internship at Microsoft., Chongyang Tao2, Tao Shen3, Can Xu2, Hongbo Xu1, Guodong Long3, Jian-guang Lou2 1Institute of Information Engineering, CAS, {xuxiaohan,hbxu}@iie.ac.cn 2Microsoft Corporation, {chotao,caxu,jlou}@microsoft.com 3AAII, School of CS, FEIT, UTS, {tao.shen,guodong.long}@uts.edu.au
@@ 1 Introduction
In the ever-evolving landscape of artificial intelligence, large language models (LLMs) have emerged as a cornerstone of natural language understanding and generation (Brown et al., 2020; Touvron et al., 2023a, a; OpenAI, 2023). As these LLMs have grown in capability, a pivotal challenge has come to the forefront: imbuing them with the ability to reason effectively. The capacity to engage in sound reasoning is a hallmark of human intelligence, enabling us to infer, deduce, and solve problems. In LLMs, this skill is paramount for enhancing their practical utility. Despite their remarkable capabilities, LLMs often struggle with nuanced reasoning (Blair-Stanek et al., 2023; Arkoudas, 2023), prompting researchers to explore innovative strategies to bolster their reasoning prowess (Wei et al., 2022b; Gao et al., 2023; Besta et al., 2023).
@@
Figure 1: Example inputs of CoT prompting versus CoT prompting with Re2. Re2 is a simple prompting method that repeats the question as input. Typically, tokens in the question, such as "tennis balls", cannot see subsequent tokens in the original setup for LLMs (the top figure). In contrast, LLMs with Re2 allows "tennis balls" in the second pass to see the entire question containing "How many …", achieving an effect of a "bidirectional" understanding (the bottom figure).
@@
Figure 2: Illustration of the attention distribution of each token in the second pass attending to the entire input in LLaMA-2 using the GSM8K dataset. A darker cell indicates higher attention. The region within the dashed upper triangle demonstrates that every token in the second pass has obvious attention to its subsequent tokens in the first pass. This suggests that re-reading in LLMs is promising for achieving a "bidirectional" understanding of the question.
@@
Existing research on reasoning has predominantly concentrated on designing diverse thought-eliciting prompting strategies to elicit reasoning processes in the output phase, such as Chain-of-Thought (CoT) Wei et al. (2022b), Program-Aided Language Model (PAL) Gao et al. (2023), etc. Yao et al. (2023a); Besta et al. (2023); Wang et al. (2023a). In contrast, scant attention has been paid to the understanding of the input phase. In fact, comprehension is the first step before solving the problem, which is crucially important. However, in the era of generative AI, most LLMs adopt the decoder-only LLMs with unidirectional attention, like GPT-3 Brown et al. (2020) and LLaMA Touvron et al. (2023b). Compared with encoder-based language models that feature bidirectional attention Devlin et al. (2019), the unidirectional attention limits every token’s visibility to only previous tokens when encoding a question, potentially impairing the global understanding of the question Du et al. (2022) (the top figure in Figure 1). Fortunately, many cognitive science studies have revealed that humans tend to re-read questions during learning and problem-solving to enhance the comprehension process Dowhower (1987, 1989); Ozek and Civelek (2006). Motivated by this, we also conduct a preliminary empirical study for LLaMA-2 Touvron et al. (2023b) by repeating the question two times using the GSM8K dataset Cobbe et al. (2021). Figure 2 shows that LLaMA-2 with re-reading is promising to achieve a "bidirectional" understanding of the question, and further improve the reasoning performance.
@@
Based on the observation and inspired by the human strategy of re-reading, we present a simple yet effective and general reasoning prompting strategy, Re2, i.e., Re-Reading the question as input (see the illustration in Figure 1). Similar to the human problem-solving process, where the primary task is to comprehend the problem, we place our emphasis on designing the prompting strategy on the input phase. Thus, Re2 is orthogonal to and compatible with most thought-eliciting prompting methods in the output phase, such as CoT, PAL, etc. Moreover, instead of processing the input in only a single pass, the repetition of questions enables LLMs to allocate more computational resources to input encoding, akin to "horizontally" increasing the depth of neural networks. Therefore, LLMs with Re2 is promising to have a deeper understanding of the question and improve reasoning performance. More interestingly, LLMs with Re2 show potential for a "bidirectional" understanding of questions in the context of unidirectional LLMs. This is because every token in the second pass could also attend to its subsequent tokens in the first pass (see illustration in Figure 1 and Figure 2).
@@ 3 Experiments
ChatGPT Vanilla 77.79 81.50 87.00 63.39 97.83 95.28 92.41
Vanilla+Re2 79.45 $\uparrow$ 1.66 84.20 $\uparrow$ 2.70 88.40 $\uparrow$ 1.40 58.27 $\downarrow$ 5.12 96.67 $\downarrow$ 1.16 94.49 $\downarrow$ 0.79 91.65 $\downarrow$ 0.76
CoT 78.77 78.70 85.60 55.91 95.50 93.70 88.61
CoT+Re2 80.59 $\uparrow$ 1.82 80.00 $\uparrow$ 1.30 86.00 $\uparrow$ 0.40 59.06 $\uparrow$ 3.15 96.50 $\uparrow$ 1.00 95.28 $\uparrow$ 1.58 89.87 $\uparrow$ 1.26
@@ 3.2 Language Models and Implementations
Re2 Prompting. We incorporate our Re2 strategy into these baseline methods to assess its impact, denoted as Vanilla+Re2 and CoT+Re2. To avoid the impact of randomness introduced by the demonstrations in a few-shot setting, we mainly assess our method in a zero-shot setting, following Chen et al. (2023); Wang et al. (2023a); Du et al. (2023). Additionally, for different tasks, we design answer-format instructions in prompts to regulate the structure of the final answer, facilitating precise answer extraction. Detailed information regarding the baseline prompting, Re2 prompting, and answer-format instructions can be found in the paper’s Appendix B.
Implementations. Our decoding strategy uses greedy decoding with a temperature setting of 0, thus leading to deterministic outputs. For these experiments, we employ two powerful backbones: ChatGPT (gpt-3.5-turbo-0613) OpenAI (2022) and davinci-003 (text-davinci-003) [^2] https://platform.openai.com/docs/models/gpt-3-5 , across all prompting methods, including Vanilla, CoT, Vanilla+Re2, and CoT+Re2.
@@ 3.3 Evaluation Results
When applied to ChatGPT, Re2 exhibits consistent improvement on most datasets, except for a slight drop in performance on a few datasets, e.g., AQUA and MultiArith, when using Vanilla+Re2. This exception could be due to ChatGPT’s exposure to these datasets with CoT outputs during instruction fine-tuning (IFT) Chen et al. (2023). On such datasets, ChatGPT with Vanilla still produces CoT-like output (see examples in Appendix E) and even outperforms ChatGPT with CoT (as indicated by the results in Tables 1 and 2). Chen et al. (2023) obtained similar experimental results and suggested that this occurs because ChatGPT may have been exposed to these task datasets containing CoT explanations without explicit prompting. Therefore, additional explicit instructions, like CoT or Re2, might disrupt this learned pattern in ChatGPT, possibly leading to decreased performance. Nonetheless, on some datasets like SVAMP, ASDIV, CSQA, and Date, Re2 still manages to improve the baseline Vanilla prompting. Moreover, in datasets where CoT prompting normally surpasses Vanilla prompting, such as GSM, StrategyQA, and Coin, Re2 significantly enhances Vanilla prompting ($\uparrow$ 4.63 on StrategyQA and $\uparrow$ 5.20 on the Coin dataset). Overall, our Re2 method still achieves improvements in 71% of the experiments on ChatGPT. More examples from the experiment results can be found in Appendix E.
@@ 3.4 Discussions
We delve deeper into the impact of the times of question re-reading on reasoning performance. Figure 3 illustrates how the performance of two distinct LLMs evolves concerning various times of question re-reading. An overarching pattern emerges across all models: performance improves until the number of re-reads reaches 2 or 3, after which it begins to decline with further increases in question re-reading times. The potential reasons for inferior performance when reading the question multiple times are two-fold: i) overly repeating questions may act as demonstrations to encourage LLMs to repeat the question rather than generate the answer, and ii) repeating the question significantly increase the inconsistency of the LLMs between our inference and pretraining/alignment (intuitively in the learning corpora, we usually repeat a question twice to emphasize the key part, rather not more). It’s noteworthy that reading the question two times tends to be optimal for accommodating most scenarios in our experiments, which is why we refer to this practice as “re-reading” in our paper.
@@
Compared to previous methods attempting to elicit thoughts in the output from LLMs, our Re2 emphasizes the understanding of the input. Therefore, we are intrigued to explore whether Re2 is effective with various thought-eliciting prompting strategies other than CoT. To investigate this, we apply Re2 to two other recently introduced prompting methods, namely, Plan-and-Solve (PS) Wang et al. (2023a) and Program-Aided Language models (PAL) Gao et al. (2023). The former model devises a plan to divide the entire task into smaller subtasks, and then carries out the subtasks according to the plan, while the latter generates programs as the intermediate reasoning steps. We directly apply our Re2 to these two methods by making a simple alteration to the input by repeating the question. Table  3 presents the evaluation findings on the GSM benchmark. Our observations reveal a consistent trend, akin to what was observed with chain-of-thought prompting. These results suggest that the effectiveness of our Re2 mechanism generally extends across various prompting methodologies.
@@
Table 6: Evaluation results of re-reading with self-consistency on GSM benchmark.
@@
Existing research indicates that the chain-of-thought prompting approach can be enhanced by adopting the self-consistency method, which involves aggregating the majority final answer from multiple sampled generations. We are also intrigued by the potential for further enhancing the proposed re-reading mechanism using this method. Consequently, we conduct experiments testing the integration of Re2 with the self-consistency approach on the GSM benchmark with temperature of 0.7 when decoding (Table 6). Our findings demonstrate that self-consistency significantly enhances the performance of both prompting methods. Despite self-consistency’s aggregation of multiple answers, our re-reading mechanism still contributes to improvement on most scenarios, indicating its compatibility with the self-consistency approach.
@@
We further investigate the impact of input question complexity on the reasoning performance of both CoT and CoT+Re2 promptings using ChatGPT on GSM8K dataset, as shown in the left part of Figure 4. In accordance with Fu et al. (2022), we measure question complexity by counting the reasoning steps present in the ground-truth explanations. Our findings reveal a noticeable trend: the performance of all promptings generally diminishes as question complexity increases, suggesting that the current LLMs still struggle with handling intricate queries. Notably, the introduction of re-reading enhances performance on various complexities, including those slightly complex questions. This observation underscores the benefits of Re2 for improving reasoning capabilities over complex questions. To further validate the improved understanding ability, we calculate the coverage degree (n-gram recall) between the generations and the input questions, as illustrated in the right part of Figure 4. The results indicate that Re2 increases the n-gram (n=1,2,3,4) recall in the output explanations, underscoring how our method enhances the model’s focus on the question during the reasoning process.
@@ 4 Related Work
LLMs represent a significant milestone in the journey towards artificial general intelligence (AGI) OpenAI (2023); Touvron et al. (2023b) . Reasoning ability is particularly crucial on the way towards AGI, where artificial intelligence needs to act or think like human beings Qiao et al. (2023); Huang and Chang (2023). In the literature on LLMs, performing reasoning tasks via interaction in natural language plays a significant role in evaluating an LLM, into which academia and industry have been dedicating many endeavors Wei et al. (2022a); Suzgun et al. (2023b); Turpin et al. (2023). In principle, most works for reasoning with large language models could fall into the paradigm of “Chain-of-Though” Wei et al. (2022b); Kojima et al. (2022), which assists LLMs in fulfilling complex reasoning tasks by generating intermediate steps explicitly. Therefore, most of the endeavors are dedicated to improving the basic principle by the following aspects: i) the structure of “chain”, e.g., tree Yao et al. (2023a), graph Yao et al. (2023b); ii) the modality of the chain, e.g., program Gao et al. (2023); iii) the reliability of the chain, e.g., self-consistency Wang et al. (2023c), faithful Lyu et al. (2023), retrieval-based verifying He et al. (2023); and iv) decomposition of the chain, e.g., least-to-most Zhou et al. (2023), decomposed Radhakrishnan et al. (2023), plan-to-solve Wang et al. (2023a). In contrast, our simple re-reading strategy for LLMs is orthogonal to these improvements via a trade-off between the intermediate steps and the query itself. Besides, our re-reading strategy is complementary to many previous works by preventing the answer from being derived overwhelmingly from the CoT but overlooking the original query.
Re-reading Strategy in Text Understanding.
@@
In deep learning, the success of performing text-understanding tasks Song et al. (2018); Luo et al. (2019a); Yang et al. (2019); Lei et al. (2019) depends on the heuristics of human reading strategy, e.g., pre-reading, re-reading and post-reading Saricoban (2002); Toprak and ALMACIOĞLU (2009); Pressley and Afflerbach (2012); Ozek and Civelek (2006); Dowhower (1989). Specifically, many effective algorithms have been crafted around the idea of re-reading. Although deep architectures, from multi-layer Bi-LSTM Huang et al. (2015) to Transformer-encoder Vaswani et al. (2017), have their mechanisms that provide a form of “re-reading”, the notion that simply processing an input once might not be sufficient for understanding or generating a complex output has been long-standing. Initially, Sha et al. (2016) and Sha et al. (2017) found that repeated reading mechanisms do improve performance on some tasks, e.g., sentiment analysis, semantic relation classification, and event extraction. Then, Liu and Li (2016) propose to mimic the repeated reading strategy and present neural networks with multi-level attention, which is proven effective in recognizing implicit discourse relations. Sequentially, Zhu et al. (2018) propose a multi-glance mechanism, modeling the habit of reading behavior, which can benefit a wide range of tasks. More recently, Luo et al. (2019b) adopt a network to encode the gist of paragraphs for rough reading and a decision-making policy for careful reading, which can improve extractive summarization. Therefore, it is natural to introduce a re-reading strategy to LLMs since the Transformer-decoder architecture of LLMs, with unidirectional attention mechanisms, hinders the implicit re-reading capability.

### v4

@@ Abstract
To enhance the reasoning capabilities of off-the-shelf Large Language Models (LLMs), we introduce a simple, yet general and effective prompting method, Re2, i.e., Re-Reading the question as input. Unlike most thought-eliciting prompting methods, such as Chain-of-Thought (CoT), which aim to elicit the reasoning process in the output, Re2 shifts the focus to the input by processing questions twice, thereby enhancing the understanding process. Consequently, Re2 demonstrates strong generality and compatibility with most thought-eliciting prompting methods, including CoT. Crucially, Re2 facilitates a "bidirectional" encoding in unidirectional decoder-only LLMs because the first pass could provide global information for the second pass. We begin with a preliminary empirical study as the foundation of Re2, illustrating its potential to enable "bidirectional" attention mechanisms. We then evaluate Re2 on extensive reasoning benchmarks across 14 datasets, spanning 112 experiments, to validate its effectiveness and generality. Our findings indicate that, with the exception of a few scenarios on vanilla ChatGPT, Re2 consistently enhances the reasoning performance of LLMs through a simple re-reading strategy. Further analyses reveal Re2’s adaptability, showing how it can be effectively integrated with different LLMs, thought-eliciting prompting, and ensemble strategies. [^1] Our code is available at Github.
@@
[…]
@@ 1 Introduction
In the ever-evolving landscape of artificial intelligence, large language models (LLMs) have emerged as a keystone of natural language understanding and generation (Brown et al., 2020; Touvron et al., 2023a, a; OpenAI, 2023; Xu et al., 2024). As LLMs have become more advanced, a key challenge has emerged: teaching them to reason effectively. The ability to reason well is a key aspect of human intelligence, allowing us to infer, deduce, and solve problems. In LLMs, this skill is crucial for improving their practical use. Despite their impressive abilities, LLMs often have difficulty with reasoning tasks (Blair-Stanek et al., 2023; Arkoudas, 2023), urging researchers to explore more strategies to bolster reasoning ability (Wei et al., 2022b; Gao et al., 2023; Besta et al., 2023).
@@
Figure 1: Example inputs of CoT prompting versus CoT prompting with Re2. In original CoT, every token in the question cannot see its later tokens since most LLMs are autoregressive models (the top figure). Re2 is a simple prompting method that repeats the question as input. LLMs with Re2 allows each token in the second pass, e.g. "tennis balls", to see its later tokens from the first pass, e.g. "How many …", achieving an effect of a "bidirectional" understanding (the bottom figure).
@@
Figure 2: Illustration of the attention distribution in LLaMA-2 by repeating the question as the input (a darker cell indicates higher attention). The region within the red dashed upper triangle demonstrates that every token in the second pass has obvious attention to its later tokens in the first pass. This suggests that re-reading in LLMs is promising for achieving a "bidirectional" understanding of the question.
@@
Existing research on reasoning has predominantly concentrated on designing diverse thought-eliciting prompting strategies to elicit reasoning processes in the output phase, such as Chain-of-Thought (CoT) Wei et al. (2022b), Program-Aided Language Model (PAL) Gao et al. (2023), etc. Yao et al. (2023a); Besta et al. (2023); Wang et al. (2023a). In contrast, scant attention has been paid to the understanding of the input phase. In fact, comprehension is the first step before solving the problem, which is crucially important. However, in the era of generative AI, most LLMs adopt the decoder-only LLMs with unidirectional attention, like GPT-3 Brown et al. (2020) and LLaMA Touvron et al. (2023b). This unidirectional attention limits every token’s visibility to only previous tokens when encoding a question, potentially impairing the bidirectional understanding of each word in the question Du et al. (2022). In Figure 1, the last sentence, "How many …", highlights the question’s main focus, which is crucial for the understanding of the preceding words. However, LLMs cannot see the subsequent words when encoding a token due to their unidirectional vision.
Fortunately, many cognitive science studies have revealed that humans tend to re-read questions during learning and problem-solving to enhance the comprehension process Dowhower (1987, 1989); Ozek and Civelek (2006). The first reading provides an overall understanding, which benefits the second reading. Motivated by this, we also conduct a preliminary empirical study for LLaMA-2 Touvron et al. (2023b) by repeating the question two times as the input using the GSM8K dataset Cobbe et al. (2021). The attention heatmap in Figure 2 shows that the re-reading strategy allows LLaMA-2 to achieve a "bidirectional" understanding of the question, which is expected to further improve the reasoning performance.
@@
Based on the observation and inspired by the human strategy of re-reading, we present a simple yet effective and general reasoning prompting strategy, Re2, i.e., Re-Reading the question as input (see the illustration in Figure 1). While our Re2 is simple, it offers several advantages for LLMs’ reasoning scenarios. (1) This approach mirrors the human strategy of problem-solving. LLMs with Re2 show potential for a "bidirectional" understanding of questions. (2) Repeating questions allows LLMs to allocate more computational resources to input encoding, similar to "horizontally" increasing the depth of neural networks. (3) Re2 emphasizes understanding during the input phase, making it orthogonal to and compatible with most thought-eliciting prompting methods that focus on the output phase, such as CoT and PAL.
@@ 3 Experiments
ChatGPT Vanilla 77.79 81.50 87.00 63.78 97.83 95.28 92.41
Vanilla+Re2 79.45 $\uparrow$ 1.66 84.20 $\uparrow$ 2.70 88.40 $\uparrow$ 1.40 59.45 $\downarrow$ 4.33 96.67 $\downarrow$ 1.16 94.49 $\downarrow$ 0.79 91.65 $\downarrow$ 0.76
CoT 78.77 78.70 85.60 55.51 95.50 93.70 88.61
CoT+Re2 80.59 $\uparrow$ 1.82 80.00 $\uparrow$ 1.30 86.00 $\uparrow$ 0.40 59.45 $\uparrow$ 3.94 96.50 $\uparrow$ 1.00 95.28 $\uparrow$ 1.58 89.87 $\uparrow$ 1.26
@@ 3.2 Language Models and Implementations
Re2 Prompting. We incorporate Re2 into these baseline methods to assess its impact, denoted as Vanilla+Re2 and CoT+Re2. To avoid the impact of randomness introduced by the demonstrations in a few-shot setting, we mainly assess our method in a zero-shot setting, following Chen et al. (2023); Wang et al. (2023a); Du et al. (2023). Additionally, for different tasks, we design answer-format instructions in prompts to regulate the format of the final answer, facilitating precise answer extraction. Detailed information regarding the baseline prompting, Re2 prompting, and answer-format instructions can be found in the Appendix B.
Implementations. Our decoding strategy uses greedy decoding with a temperature setting of 0, thus leading to deterministic outputs. For these experiments, we employ two powerful backbones: ChatGPT (gpt-3.5-turbo-0613) OpenAI (2022) and davinci-003 (text-davinci-003) [^2] https://platform.openai.com/docs/models/gpt-3-5 , across all prompting methods, including Vanilla, CoT, Vanilla+Re2, and CoT+Re2. We also test Re2 on more advanced GPT-4o-mini in Appendix C.
@@ 3.3 Evaluation Results
When applied to ChatGPT, Re2 exhibits consistent improvement on most datasets, except for a slight drop in performance on a few datasets, e.g., AQUA and MultiArith, when using Vanilla+Re2. This exception could be due to ChatGPT’s exposure to these datasets with CoT outputs during instruction fine-tuning (IFT) Chen et al. (2023). On such datasets, ChatGPT with Vanilla still produces CoT-like output (see examples in Appendix G) and even outperforms ChatGPT with CoT (as indicated by the results in Tables 1 and 2). Chen et al. (2023) obtained similar experimental results and suggested that this occurs because ChatGPT may have been exposed to these task datasets containing CoT explanations without explicit prompting. Therefore, additional explicit instructions, like CoT or Re2, might disrupt this learned pattern in ChatGPT, possibly leading to decreased performance. Nonetheless, on some datasets like SVAMP, ASDIV, CSQA, and Date, Re2 still manages to improve the baseline Vanilla prompting. Moreover, in datasets where CoT prompting normally surpasses Vanilla prompting, such as GSM, StrategyQA, and Coin, Re2 significantly enhances Vanilla prompting ($\uparrow$ 4.63 on StrategyQA and $\uparrow$ 5.20 on the Coin dataset). Overall, our Re2 method still achieves improvements in 71% of the experiments on ChatGPT. More examples from the experiment results can be found in Appendix G.
@@ 3.4 Discussions
We delve deeper into the impact of the times of question re-reading on reasoning performance. Figure 3 illustrates how the performance of two distinct LLMs evolves concerning various times of question re-reading. An overarching pattern emerges across all models: performance improves until the number of re-reads reaches 2 or 3, after which it begins to decline with further increases in question re-reading times. The potential reasons for inferior performance when reading the question multiple times are two-fold: i) appropriate reading times increase LLMs’ ability to generate correct answers. However, excessively repeating questions may serve as demonstrations, causing the LLMs to repeat the questions themselves (see Appendix E for detailed analysis). and ii) repeating the question significantly increase the inconsistency of the LLMs between our inference and pretraining/alignment (intuitively in the learning corpora, we usually repeat a question twice). It’s noteworthy that reading the question twice is optimal in most scenarios, which is why we refer to it as “re-reading” in our paper.
@@
Compared to previous methods attempting to elicit thoughts in the output from LLMs, our Re2 emphasizes the understanding of the input. Therefore, we are intrigued to explore whether Re2 is effective with various thought-eliciting prompting strategies other than CoT. To investigate this, we apply Re2 to two other recently introduced prompting methods, namely, Plan-and-Solve (PS) Wang et al. (2023a) and Program-Aided Language models (PAL) Gao et al. (2023). The former model devises a plan to divide the entire task into smaller subtasks, and then carries out the subtasks according to the plan, while the latter generates programs as the intermediate reasoning steps. We directly apply our Re2 to these two methods by making a simple alteration to the input by repeating the question. Table  3 presents the evaluation findings on the GSM benchmark. Our observations reveal a consistent trend, akin to what was observed with CoT prompting. These results suggest that the effectiveness of our Re2 generally extends across various prompting methodologies.
@@
Table 6: Evaluation results of re-reading with self-consistency (t-test, $\dagger$ denote p-value < 0.05).
@@
Existing research indicates that the chain-of-thought prompting approach can be enhanced by adopting the self-consistency method, which involves aggregating the majority final answer from multiple sampled generations. We are also intrigued by the potential for further enhancing the proposed re-reading mechanism using this method. Consequently, we conduct experiments testing the integration of Re2 with the self-consistency approach on the GSM benchmark by using ChatGPT. The temperature is set to 0.7. We report the results averaged over 10 runs, where we sampled 10 outputs independently from the LLMs in each run. Table 6 demonstrate that self-consistency significantly enhances the performance of both prompting methods. Despite self-consistency’s aggregation of multiple answers, our re-reading mechanism still contributes to improvement on most scenarios, indicating its compatibility with the self-consistency approach.
@@
We further investigate the impact of input question complexity on the reasoning performance of both CoT and CoT+Re2 promptings using ChatGPT on GSM8K dataset, as shown in the left part of Figure 4. In accordance with Fu et al. (2022), we measure question complexity by counting the reasoning steps present in the ground-truth explanations. Our findings reveal that the performance of all promptings generally diminishes as question complexity increases, suggesting that the current LLMs still struggle with handling intricate queries. Notably, the introduction of re-reading enhances performance on various complexities, including those slightly complex questions. This observation underscores the benefits of Re2 for improving reasoning capabilities over complex questions. To further validate the improved understanding ability, we calculate the coverage degree (n-gram recall) between the generations and the input questions, as illustrated in the right part of Figure 4. The results indicate that Re2 increases the n-gram (n=1,2,3,4) recall in the output explanations, underscoring how our method enhances the model’s focus on the question during the reasoning process.
@@
Figure 5: Re2’s impact on inference efficiency and GPU memory usage.
Impact on Inference Efficiency and Memory Usage.
Re2 doubles the question length in both zero- and few-shot settings, which may affect inference efficiency and memory usage. This section quantitatively explores that impact. We utilize Llama-2 7B with float16 precision and randomly sample 100 instances from the GSM8K dataset. We measure the average inference time and memory usage across four scenarios: Zero-shot, Zero-shot + CoT, Few-shot, and Few-shot + CoT. When applying Re2, the questions in the demonstrations are also repeated. All experiments are performed on 8$\times$×NVIDIA GeForce RTX 4090 GPUs, with results shown in Figure 5. The findings reveal that RE2 only marginally increases inference time and memory usage in both zero-shot and few-shot settings, even with longer inputs. This minimal impact is attributed to various optimization and inference acceleration techniques in current LLMs, such as grouped-query attention Touvron et al. (2023b), CUDA, and GPU-based computations. For instance, grouped-query attention is particularly advantageous for long inputs, significantly accelerating decoder inference. Likewise, CUDA and GPU-based computations are highly optimized for parallel processing, especially for matrix multiplications in LLMs NVIDIA (2024).
@@ 4 Related Work
LLMs represent a significant milestone in the journey towards artificial general intelligence (AGI) OpenAI (2023); Touvron et al. (2023b) . Reasoning ability is particularly crucial on the way towards AGI, where artificial intelligence needs to act or think like human beings Qiao et al. (2023); Huang and Chang (2023). In the literature on LLMs, performing reasoning tasks via interaction in natural language plays a significant role in evaluating an LLM, into which academia and industry have been dedicating many endeavors Wei et al. (2022a); Suzgun et al. (2023b); Turpin et al. (2023). In principle, most works for reasoning with large language models could fall into the paradigm of “Chain-of-Thought” Wei et al. (2022b); Kojima et al. (2022), which assists LLMs in fulfilling complex reasoning tasks by generating intermediate steps explicitly. Therefore, most of the endeavors are dedicated to improving the basic principle by the following aspects: i) the structure of “chain”, e.g., tree Yao et al. (2023a), graph Yao et al. (2023b); ii) the modality of the chain, e.g., program Gao et al. (2023); iii) the reliability of the chain, e.g., self-consistency Wang et al. (2023c), faithful Lyu et al. (2023), retrieval-based verifying He et al. (2023); and iv) decomposition of the chain, e.g., least-to-most Zhou et al. (2023), decomposed Radhakrishnan et al. (2023), plan-to-solve Wang et al. (2023a). In contrast, our simple re-reading strategy for LLMs is orthogonal to these improvements via a trade-off between the intermediate steps and the query itself. Besides, our re-reading strategy is complementary to many previous works by preventing the answer from being derived overwhelmingly from the CoT but overlooking the original query.
Re-reading Strategy in NLP.
@@
In deep learning, the success of performing text-understanding tasks Song et al. (2018); Luo et al. (2019a); Yang et al. (2019); Lei et al. (2019) depends on the heuristics of human reading strategy, e.g., pre-reading, re-reading and post-reading Saricoban (2002); Toprak and ALMACIOĞLU (2009); Pressley and Afflerbach (2012); Ozek and Civelek (2006); Dowhower (1989). Specifically, many effective algorithms have been crafted around the idea of re-reading. Although deep architectures, from multi-layer Bi-LSTM Huang et al. (2015) to Transformer-encoder Vaswani et al. (2017), have their mechanisms that provide a form of “re-reading”, the notion that simply processing an input once might not be sufficient for understanding or generating a complex output has been long-standing. Initially, Sha et al. (2016) and Sha et al. (2017) found that repeated reading mechanisms do improve performance on some tasks, e.g., sentiment analysis, semantic relation classification, and event extraction. Then, Liu and Li (2016) propose to mimic the repeated reading strategy and present neural networks with multi-level attention, which is proven effective in recognizing implicit discourse relations. Sequentially, Zhu et al. (2018) propose a multi-glance mechanism, modeling the habit of reading behavior, which can benefit a wide range of tasks. Luo et al. (2019b) adopt a network to encode the gist of paragraphs for rough reading and a decision-making policy for careful reading, which can improve extractive summarization. More recently, Springer et al. (2024) have shown the effectiveness of repeating input to get bidirectional embeddings on text embedding tasks. Therefore, it is natural to introduce a re-reading strategy to LLMs’ reasoning, since the Transformer-decoder architecture of LLMs with unidirectional attention mechanisms hinders the implicit bidirectional capability.